OpenStudy (anonymous):

You pick 5 cards out of a deck of 52. What is the probability that you get exactly two spades?

OpenStudy (anonymous):

I tried it and thought it was:\[\left(\begin{matrix}5 \\ 2\end{matrix}\right) (\frac{ 1 }{ 4 })^2 (\frac{ 3 }{ 4 })^3\] But someone else posted somewhere on the internet that it would be \[\frac{ 13 }{ 52 } * \frac{ 12 }{ 51 } * \frac{ 39 }{ 50 }* \frac{ 38 }{ 49 } * \frac{ 37 }{ 48 }\] It doesn't say anything about replacement of the cards